# CWTHF: Identifying Dark Matter Halos with Continuous Wavelet Transform

### v1

@@ 1 Introduction
In such simulations, the density of the universe is typically represented by particles, including dark matter, gas, and stars. To extract structural information from these particles, it is essential to identify isolated, gravitationally bound objects, namely dark matter halos. In recent years, the rapid advancement of computer science has enabled researchers to conduct simulations with billions or even trillions of particles. This increase in data volume necessitates the development of faster algorithms for data processing. Furthermore, while it is commonly stated that a dark matter halo is a ‘gravitationally bound object’, this alone is insufficient to definitively identify halos from simulation datasets (Knebe et al., 2013). The precise definition of a dark matter halo remains a contentious issue. The lack of a universally accepted definition, coupled with the need to analyze simulation data, has driven the development of halo-finding algorithms. Below, we introduce some of them roughly according to the timeline.
@@
The halo finders such as HOP and SKID are designed using the kernel density estimation (KDE) method. While the KDE method is capable of deriving a continuous density distribution from discrete particle data (Li & He, 2018), the wavelet method, by contrast, not only can obtain a continuous density distribution from discrete particle data, but also offers a distinct halo boundary. In a previous study, we employed the continuous wavelet transform (CWT) technique to identify halos from a pseudo-two-dimensional (2D) dataset, which was derived from three-dimensional (3D) simulation data through a process of dimensional reduction (Li et al., 2024). Our work has demonstrated the applicability of the CWT in identifying halos from 2D dataset. The wavelet transform technique has become an important research method in cosmology, utilized for extracting information about the clustering of cosmic structures, owing to its ability to analyze data in both spatial and scale domains simultaneously (e.g., Freeman et al., 2002; Vavilova & Babyk, 2018; Wang et al., 2008; Ciprini et al., 2007; Djafer et al., 2012; González-Gómez et al., 2010; Patrikeev et al., 2006). The structures present in the given data will result in a high CWT value at the location of these structures, and the size of these structures can be determined by the scale of the cross-scale maximum (Bendjoya et al., 1991; Bijaoui et al., 1992; Slezak et al., 1993; Grebenev et al., 1995; Flin & Krywult, 2006). Furthermore, the precise configuration of these structures can be ensured by segmenting the CWT from local maxima (Flin & Krywult, 2006; Baluev & Rodionov, 2020).
The current paper presents an updated and optimized version of our previous 2D algorithm to the real 3D case, introducing the continuous wavelet transform halo finder (CWTHF). Unlike traditional methods that directly identify structures in the density field, the CWTHF employs a multi-scale analysis to identify halos in the 4D (3D physical plus 1D scale) wavelet space. This methodology allows for the selection of specific scales for investigation, thereby enabling the identification of large halos with minimal resources in a time-efficient manner. It is important to note that any halo finder inherently has its own definition of a halo. In the case of CWTHF, the position and extent of the halos are contingent upon the location of the cross-scale maxima and the declining CWT value.
@@ 2.4 Cross-scale Verification and Density Threshold
The threshold density is considerably lower than that predicted by the spherical collapse model. This discrepancy can be attributed to two factors. First, the program uses the grid volume instead of the virial volume to quantify the volume of these structures, as shown in Figure 5. In the case of low-to-medium mass structures, the grid volume is larger than the virial volume because it already includes the low-density external region. Second, regarding the density criterion, the goal is simply to filter out suspicious maxima, rather than to achieve the final identification result. During the density check phase, the structures input to the program are only a rough segmentation of these particles. Applying a strict criterion at this stage could potentially introduce bias into the final outcome.
@@ Acknowledgments
MXL thanks anonymous users on Github for their assistance with some coding techniques. The authors also express their gratitude to the SIMBA team for making their data publicly available. The source code for CWTHF is available on GitHub at https://github.com/salty-fish-114514/CWTHF/tree/main.

### v2

@@ 1 Introduction
In such simulations, the density of the universe is typically represented by particles, including dark matter, gas, and stars. To extract structural information from these particles, it is essential to identify isolated, gravitationally bound objects, namely dark matter halos. In recent years, the rapid advancement of computer science has enabled researchers to conduct simulations with billions or even trillions of particles (e.g., Davé et al., 2019; Nelson et al., 2019; Villaescusa-Navarro et al., 2020). Although some researchers can utilize these data without identifying isolated halos (e.g., Balaudo et al., 2024), the increasing volume of data necessitates the development of faster algorithms for data processing. Furthermore, while it is commonly stated that a dark matter halo is a ‘gravitationally bound object’, this alone is insufficient to definitively identify halos from simulation datasets (Knebe et al., 2013). The precise definition of a dark matter halo remains a contentious issue. The lack of a universally accepted definition, coupled with the need to analyze simulation data, has driven the development of halo-finding algorithms. Below, we introduce some of them roughly according to the timeline.
@@
The halo finders, such as HOP and SKID, are designed using the kernel density estimation (KDE) method. While the KDE method is capable of deriving a continuous density distribution from discrete particle data (Li & He, 2018), the wavelet method, by contrast, not only can obtain a continuous density distribution from discrete particle data, but also offers a distinct halo boundary. In a previous study, we employed the continuous wavelet transform (CWT) technique to identify halos from a pseudo-two-dimensional (2D) dataset, which was derived from three-dimensional (3D) simulation data through a process of dimensional reduction (Li et al., 2024). Our work has demonstrated the applicability of the CWT in identifying halos from a 2D dataset. The wavelet transform technique has become a pivotal tool in cosmology, enabling the analysis of cosmic structure clustering across both spatial and scale domains (e.g., Freeman et al., 2002; Patrikeev et al., 2006; Ciprini et al., 2007; Romeo et al., 2008; Wang et al., 2008; González-Gómez et al., 2010; Djafer et al., 2012; Vavilova & Babyk, 2018; Wang et al., 2022; Wang & He, 2022). The structures present in the given data will result in a high CWT value at the location of these structures, and the size of these structures can be determined by the scale of the cross-scale maximum (Bendjoya et al., 1991; Bijaoui et al., 1992; Slezak et al., 1993; Grebenev et al., 1995; Flin & Krywult, 2006). Furthermore, the precise configuration of these structures can be ensured by segmenting the CWT from local maxima (Flin & Krywult, 2006; Baluev & Rodionov, 2020).
The current paper presents an updated and optimized version of our previous 2D algorithm to the real 3D case, introducing the continuous wavelet transform halo finder (CWTHF). Unlike traditional methods that directly identify structures in the density field, the CWTHF employs a multi-scale analysis to identify halos in the 4D (3D physical plus 1D scale) wavelet space. This methodology allows for the selection of specific scales for investigation, thereby enabling the identification of large halos with minimal resources in a time-efficient manner. The multiscale technique is a powerful tool in cosmology and has been used to detect various cosmological features (Aragón-Calvo et al., 2007). It is important to note that any halo finder inherently has its own definition of a halo. In the case of CWTHF, the position and extent of the halos are contingent upon the location of the cross-scale maxima and the declining CWT value.
@@ 2.4 Cross-scale Verification and Density Threshold
The threshold density is considerably lower than that predicted by the spherical collapse model, any value greater than 1 and less than 7 produces acceptable results (see parametric studies presented in Section 4). This discrepancy can be attributed to two factors. First, the program uses the grid volume instead of the virial volume to quantify the volume of these structures, as shown in Figure 5. In the case of low-to-medium mass structures, the grid volume is larger than the virial volume because it already includes the low-density external region. Second, regarding the density criterion, the goal is simply to filter out suspicious maxima, rather than to achieve the final identification result. During the density check phase, the structures input to the program are only a rough segmentation of these particles. Applying a strict criterion at this stage could potentially introduce bias into the final outcome.
@@ 3 Results
Both CWT halos exhibit an elliptical central structure, confirming the ability of the spherical CWT to identify non-spherical structures. For these irregular halos, their high-density centers still align with the wavelets of similar size, thereby forming peaks in the 4D wavelet space. Furthermore, it is important to note that the information regarding the halo shape in the CIC density is not completely erased in the convolution process. The boundaries delineated in the CWT grid follow the shapes of the halos in the CIC grid, and the outward extension during the segmentation process further enriches the shapes of the CWT halos.
@@ 5 Conclusions and Discussions
Our program still has great potential for future optimizations. Even with the fast algorithm, the actual performance of the CWTHF has no obvious advantage over that of the FOF. In addition, the memory usage, which is associated with the kw_high, depends on the size of the simulation box. For simulations with larger box lengths, it is necessary to set a higher kw_high in order to access small-scale halos. This, of course, results in higher memory usage. Fortunately, an increase in box length generally corresponds to a decrease in spatial resolution. This enables us to avoid the necessity of extending the grid to a high value, thereby mitigating the problematic memory usage in these instances.
@@ Acknowledgments
The authors thank Dr. Miguel Aragon for his helpful comments and suggestions. MXL thanks anonymous users on Github for their assistance with some coding techniques. The authors also express their gratitude to the SIMBA team for making their data publicly available. The source code for CWTHF is available on Zenodo (Li, 2025)444For future updated version of the CWTHF, see https://github.com/salty-fish-114514/CWTHF/tree/main.